OpenStudy (anonymous):

(3x+5)/(4)- (2x-1)/(3)=2 solve fir x.

OpenStudy (anonymous):

\[\frac{ 3x+5 }{ 4 } - \frac{ 2x-1 }{ 3 } = 2\] \[3*(3x+5) - 4*( 2x-1 ) = 2*12\] \[9x+15 -8x+4 =24\] \[x+19=24\] \[x=5\]